Ten percent of Americans are allergic to ragweed. Suppose a random sample of 200 people is selected.

1. Identify n    

2. Identify p   (Round to 2 decimal places)

3. Identify 1 – p   (Round to 2 decimal places)


 


1
Expert's answer
2021-07-08T11:10:03-0400

1. "n=200"

2. "p= 10 \; \% = 0.10"

3. "1-p=1-0.10=0.90"